# Markov kernels in Mathlib’s probability library

### v1

@@ 2.3 Copy-discard and Markov categories
We can form several categories with measurable spaces as objects and kernels as morphisms. Those are not defined as categories in Mathlib, but every result needed to show the properties of those categories is present in the library and ready to be used if someone later needs to consider kernels as a copy-discard or Markov category. We briefly mention them here to explain how they relate to the objects presented so far. For an explanation of the different types of categories named below and the theory of Markov categories, see [CJ19, Fri20].
Measurable spaces and s-finite kernels are a copy-discard category (and in particular a symmetric monoidal category) with the following data:
@@
If we restrict the morphisms to Markov kernels, we get a Markov category, which is a copy-discard category in which for every morphism $\kappa$, $d_{\mathcal{Y}}\circ\kappa=d_{\mathcal{X}}$.
2.4 The Ionescu-Tulcea theorem
We highlight a recent application of the kernel framework: the Ionescu-Tulcea theorem, which was implemented in Lean and added to Mathlib by Etienne Marion [Mar25].
Theorem 2.1 (Ionescu-Tulcea).
Let $(X_{n})_{n\in\mathbb{N}}$ be a family of measurable spaces. Let $(\kappa_{n})_{n\in\mathbb{N}}$ be a family of Markov kernels such that for any $n$, $\kappa_{n}$ is a kernel from $\prod_{i=0}^{n}X_{i}$ to $X_{n+1}$. Then there exists a unique Markov kernel $\xi:X_{0}\rightsquigarrow\prod_{i=1}^{\infty}X_{i}$ such that for any $n\geq 1$, $\pi_{[1,n]*}\xi=\kappa_{0}\otimes_{k}\ldots\otimes_{k}\kappa_{n-1}$. Here $\pi_{[1,n]}:\prod_{i=1}^{\infty}X_{i}\to\prod_{i=1}^{n}X_{i}$ is the projection on the first $n$ coordinates.
In the formalization of that Theorem, an important difficulty is the lack of associativity of products of types, as explained in the previous subsection about the composition-product. The article [Mar25] discusses this issue in detail and presents a formalization strategy to avoid inserting associators everywhere.
@@ 3 Disintegration and conditional distributions
We say that a kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}\times\mathcal{Z}$ is disintegrated by a kernel $\eta:\mathcal{X}\times\mathcal{Y}\rightsquigarrow\mathcal{Z}$ if $\kappa=\kappa_{\mathrm{fst}}\otimes\eta$. We also say that $\eta$ is a conditional kernel for $\kappa$. In the particular case of a product measure $\rho$ on $\mathcal{Y}\times\mathcal{Z}$ (which corresponds of the special case of $\mathcal{X}$ the unit space), disintegration means that $\rho$ can be expressed with the measure $\rho_{\mathrm{fst}}$ on $\mathcal{Y}$ and a kernel $\eta:\mathcal{Y}\rightsquigarrow\mathcal{Z}$ as $\rho=\rho_{\mathrm{fst}}\otimes\eta$: a joint measure can be decomposed into a marginal measure and a conditional distribution.
In general measurable spaces, disintegration is not always possible (refer to [Fad85] for exact hypotheses). However, it is possible for every finite kernel if we add assumptions on the spaces.
@@
A standard Borel space is a measurable space such that there exists a Polish topology on the underlying set such that the $\sigma$-algebra is the Borel $\sigma$-algebra of that topology.
We note that a standard Borel space has a countably generated $\sigma$-algebra. Furthermore, any standard Borel space is isomorphic as a measurable space to a subset of $\mathbb{R}$. Most spaces used in applied mathematics are standard Borel.
We first introduced the disintegration in Mathlib as a definition: under assumptions on the spaces, we built a kernel $\eta$ disintegrating a given kernel $\kappa$. Since then, a predicate has been introduced to express that a kernel disintegrates another.
@@
The advantage of that predicate is that it allows us to work with disintegrations in more general spaces, if we happen to have another mechanism to know that a kernel disintegrates another. A conditional kernel is almost everywhere equal to a Markov kernel, and almost everywhere unique. Futhermore, if $\eta$ disintegrates $\kappa$ and $\eta$ is $\kappa_{\mathrm{fst}}$-almost everywhere equal to $\eta^{\prime}$, then $\eta^{\prime}$ also disintegrates $\kappa$. That means that if $\kappa$ admits a disintegration, then there is a true Markov kernel that disintegrates $\kappa$.
The main result of this section is the following theorem, which states that there exists a conditional kernel under assumptions on the spaces. It is formalized in Mathlib as a definition Kernel.condKernel that gives a conditional kernel, which we ensure is Markov, together with an instance of the class IsCondKernel.
@@
As a corollary, if $\mathcal{Z}$ is a standard Borel space then every finite measure on $\mathcal{Y}\times\mathcal{Z}$ admits a disintegration. The proof goes through the construction of densities for finite kernels, and in particular it also gives a formalization of a Radon-Nikodym theorem. In order to state that theorem, we need to introduce the definition of the kernel with a density with respect to another kernel. For $f:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{+,\infty}$ a jointly measurable function and $\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$ an s-finite kernel, we define the kernel $f\cdot\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$ (Kernel.withDensity) by
$(f\cdot\eta)(x)(B)=\int_{y\in B}f(x,y)d\eta(x)\>.$
@@
Suppose that either $\mathcal{X}$ is countable or $\mathcal{Y}$ has a countably generated $\sigma$-algebra. Let $\kappa,\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$ be finite kernels. Then there exists a jointly measurable function $\frac{d\kappa}{d\eta}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{+,\infty}$ and a finite kernel $\kappa_{\perp\eta}:\mathcal{X}\rightsquigarrow\mathcal{Y}$ such that for every $x\in\mathcal{X}$, $\kappa_{\perp\eta}(x)$ is singular with respect to $\eta(x)$ and $\kappa=\frac{d\kappa}{d\eta}\cdot\eta+\kappa_{\perp\eta}$ .
@@ 3.1 Proof of the disintegration theorem
The main result from which both the disintegration theorem and the Radon-Nikodym theorem are derived is the existence of densities for finite kernels. For two kernels $\kappa,\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$, we write $\kappa\leq\eta$ if for all $x\in\mathcal{X}$ and measurable set $B$ of $\mathcal{Y}$, $\kappa(x)(B)\leq\eta(x)(B)$.
@@
Let $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}\times\mathcal{Z}$ and $\nu:\mathcal{X}\rightsquigarrow\mathcal{Y}$ be two finite kernels with $\kappa_{\mathrm{fst}}\leq\nu$, where $\mathcal{Y}$ has a countably generated $\sigma$-algebra. Then there exists a function $d_{\kappa,\nu}:\mathcal{X}\times\mathcal{Y}\to\mathcal{P}(\mathcal{Z})\to\mathbb{R}$ jointly measurable on $\mathcal{X}\times\mathcal{Y}$ such that for all $x\in\mathcal{X}$ and all measurable sets $B$ of $\mathcal{Y}$ and $C$ of $\mathcal{Z}$, $\int_{y\in B}d_{\kappa,\nu}(x,y)(C)\,d\nu(x)=\kappa(x)(B\times C)$.
We added a definition for the density, Kernel.density, and its properties can be found in the same file.
@@
Using the fact that a standard Borel space embeds measurably into $\mathbb{R}$ (embeddingReal), we can reduce the proof of Theorem 3.2 to the case where $\mathcal{Z}=\mathbb{R}$: for a general standard Borel space, we map the kernel to one with values in $\mathbb{R}$, disintegrate it, and then pull back the conditional kernel to the original space. Thus from a finite kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}\times\mathbb{R}$, we want to build a kernel $\eta:\mathcal{X}\times\mathcal{Y}\rightsquigarrow\mathbb{R}$ such that $\kappa=\kappa_{\mathrm{fst}}\otimes\eta$. Since $\eta$ takes values in the measures on $\mathbb{R}$, we can build them from their cumulative distribution functions. A cumulative distribution function (CDF) is a nondecreasing right-continuous function $F:\mathbb{R}\to\mathbb{R}$ with limit 0 at -$\infty$ and limit 1 at $\infty$, and characterizes a probability measure on $\mathbb{R}$. Nondecreasing right-continuous functions $\mathbb{R}\to\mathbb{R}$ are bundled in Mathlib in a type StieltjesFunction and the corresponding measure can be obtained with StieltjesFunction.measure.
@@
If $\mathcal{X}$ is countable, we can provide for each $x\in\mathcal{X}$ a measurable function $f:\mathcal{Y}\to\mathbb{Q}\to\mathbb{R}$ and proceed as above to obtain a kernel $\mathcal{Y}\rightsquigarrow\mathbb{R}$. Since $\mathcal{X}$ is countable, measurability is not an issue and we can put those kernel together into a kernel $\mathcal{X}\times\mathcal{Y}\rightsquigarrow\mathbb{R}$. If on the other hand $\mathcal{X}$ is not countable, we can’t proceed separately for each $x\in\mathcal{X}$ and have to build a function which is measurable on the product. We are able to do so if $\mathcal{Y}$ has a countably generated $\sigma$-algebra thanks to the density provided by Theorem 3.4. The density $d_{\kappa,\nu}:\mathcal{X}\times\mathcal{Y}\to\mathcal{P}(\mathbb{R})\to\mathbb{R}$ is jointly measurable on $\mathcal{X}\times\mathcal{Y}$ and such that for all $x\in\mathcal{X}$ and all measurable sets $B$ of $\mathcal{Y}$ and all $q\in\mathbb{Q}$, $\int_{y\in B}d_{\kappa,\nu}(x,y)((-\infty,q])\,d\nu(x)=\kappa(x)(B\times(-\infty,q])$. It thus satisfies the integral condition, and we are able to check that the other conditions are also fulfilled.
The countability of $\mathbb{Q}$ is important when proving the right-continuity. We want to prove that for almost all $p\in\mathcal{X}\times\mathcal{Y}$, $d_{\kappa,\nu}(p)$ is right-continuous on $\mathbb{Q}$ . Since $\mathbb{Q}$ is countable, it is enough to prove that for all $q\in\mathbb{Q}$, for almost all $p$, $d_{\kappa,\nu}(p)$ is right-continuous at $q$. To prove the almost everywhere equality of the limit on the right and the value of the function at a point $q\in\mathbb{Q}$, it suffices to prove that the integrals of those functions are equal on any measurable set $A$ of $\mathcal{X}\times\mathcal{Y}$. And since we know what the integrals of $d_{\kappa,\nu}$ are, we can prove that equality.
@@
From $\kappa,\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$ two finite kernels we want to build a jointly measurable function $\frac{d\kappa}{d\eta}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{+,\infty}$ and a finite kernel $\kappa_{\perp\eta}:\mathcal{X}\rightsquigarrow\mathcal{Y}$ such that for every $x\in\mathcal{X}$, $\kappa_{\perp\eta}(x)$ is singular with respect to $\eta(x)$ and $\kappa=\frac{d\kappa}{d\eta}\cdot\eta+\kappa_{\perp\eta}$ .
@@
Let $\kappa,\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$ be two finite kernels with $\kappa\leq\eta$, in which $\mathcal{Y}$ has a countably generated $\sigma$-algebra. Then there exists a function $d_{\kappa,\eta}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ jointly measurable on $\mathcal{X}\times\mathcal{Y}$ such that for all $x\in\mathcal{X}$ and all measurable sets $B$ of $\mathcal{Y}$, $\int_{y\in B}d_{\kappa,\eta}(x,y)\,d\eta(x)=\kappa(x)(B)$.
In the Radon-Nikodym Theorem 3.3, we don’t have the condition $\kappa\leq\eta$. We apply that corollary to $\kappa$ and $\kappa+\eta$, which satisfy that condition. The Radon-Nikodym derivative $\frac{d\kappa}{d\eta}$ is defined by the ratio $\frac{d_{\kappa,\kappa+\eta}}{1-d_{\kappa,\kappa+\eta}}$. It is then not hard to check that it satisfies the required properties, with $\kappa_{\perp\eta}=\kappa-\frac{d\kappa}{d\eta}\cdot\eta$. Thanks to the Radon-Nikodym Theorem 3.3 which is now proved, the condition $\kappa\leq\eta$ in Corollary 3.5 can be relaxed to absolute continuity of $\kappa$ with respect to $\eta$ (that is, absolute continuity of $\kappa(x)$ with respect to $\eta(x)$ for all $x$).
@@
We now give a sketch of the proof of Theorem 3.4. Our proof follows [Kal21, Theorem 9.27]. If we were interested in a single $x\in\mathcal{X}$, we could use the Radon-Nikodym theorem for measures to obtain a density. For $x\in\mathcal{X}$, $y\in\mathcal{Y}$ and $C$ a measurable set of $\mathcal{Z}$, let $\kappa(x)_{|\mathcal{Y}\times C}$ be the restriction of the measure $\kappa(x)$ to the set $\mathcal{Y}\times C$. Let then $(\kappa(x)_{|\mathcal{Y}\times C})_{\mathrm{fst}}$ be the marginal of that measure on $\mathcal{Y}$. Then we could define the density as $d^{\prime}_{\kappa,\nu}(x,y)(C)=\frac{d(\kappa(x)_{|\mathcal{Y}\times C})_{\mathrm{fst}}}{d\nu(x)}(y)$ and it would satisfy the integral condition and be measurable in $y$. The problem is to obtain a density that is jointly measurable on $\mathcal{X}\times\mathcal{Y}$.

### v2

@@ 2.3 Copy-discard and Markov categories
We can form several categories with measurable spaces as objects and kernels as morphisms. Our work on kernels was used by Gaëtan Serré to define copy-discard and Markov categories for kernels in Mathlib. We briefly mention them here to explain how they relate to the objects presented so far. For an explanation of the different types of categories named below and the theory of Markov categories, see [cho2019disintegration, fritz2020synthetic].
Measurable spaces and s-finite kernels are a copy-discard category (and in particular a symmetric monoidal category) SFinKer with the following data:
@@
If we restrict the morphisms to Markov kernels to form the category Stoch, we get a Markov category, which is a copy-discard category in which for every morphism $\kappa$, $_{\mathcal{Y}}\circ\kappa=!_{\mathcal{X}}$.
@@ 3 Disintegration and conditional distributions
We say that a kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}\times\mathcal{Z}$ is disintegrated by a kernel $\eta:\mathcal{X}\times\mathcal{Y}\rightsquigarrow\mathcal{Z}$ if $\kappa=\kappa_{\mathrm{fst}}\otimes_{k}\eta$. We also say that $\eta$ is a conditional kernel for $\kappa$. In the particular case of a product measure $\rho$ on $\mathcal{Y}\times\mathcal{Z}$ (which corresponds to the special case where $\mathcal{X}$ the unit space), disintegration means that $\rho$ can be expressed with the measure $\rho_{\mathrm{fst}}$ on $\mathcal{Y}$ and a kernel $\eta:\mathcal{Y}\rightsquigarrow\mathcal{Z}$ as $\rho=\rho_{\mathrm{fst}}\otimes_{m}\eta$ : a joint measure can be decomposed into a marginal measure and a conditional distribution.
In general measurable spaces, disintegration is not always possible (refer to [faden1985existence] for exact hypotheses). However, it is possible for every finite kernel if we add assumptions on the spaces.
@@
A standard Borel space is a measurable space such that there exists a Polish topology (second-countable topology that can be endowed with a metric for which it is complete) on the underlying set such that the $\sigma$-algebra is the Borel $\sigma$-algebra of that topology (the $\sigma$-algebra generated by the open sets).
We note that a standard Borel space has a countably generated $\sigma$-algebra. Furthermore, any standard Borel space embeds measurably into $\mathbb{R}$ (it is isomorphic as a measurable space to a subset of $\mathbb{R}$). Most spaces used in applied mathematics are standard Borel.
We first introduced the disintegration in Mathlib as a definition: under assumptions on the spaces, we built a kernel $\eta$ disintegrating a given kernel $\kappa$. Since then, Yaël Dillies and Kalle Kytölä introduced a predicate to express that a kernel disintegrates another.
@@
The command class makes this structure inferable by the typeclass system of Lean. The protected modifier means that even if the namespace IsCondKernel is open, the field disintegrate can only be accessed with the full name IsCondKernel.disintegrate, which is useful to avoid name clashes with another lemma named disintegrate in the library.
The advantage of that predicate is that it allows us to work with disintegrations in more general spaces, if we happen to have another mechanism to know that a kernel disintegrates another. A conditional kernel is almost everywhere equal to a Markov kernel, and almost everywhere unique. Futhermore, if $\eta$ disintegrates $\kappa$ and $\eta$ is $\kappa_{\mathrm{fst}}$-almost everywhere equal to $\eta^{\prime}$, then $\eta^{\prime}$ also disintegrates $\kappa$. That means that if $\kappa$ admits a disintegration, then there is a true Markov kernel that disintegrates $\kappa$. All those facts are formalized in Mathlib.
The main result of this section is the following theorem, which states that there exists a conditional kernel under assumptions on the spaces.
@@
It is formalized in Mathlib as a definition Kernel.condKernel that gives a conditional kernel (constructed following the steps outlined in Section 3.1), which we ensure is Markov, together with an instance of the class IsCondKernel.
[…]
κ.IsCondKernel κ.condKernel
Once this is done, the particular construction of the conditional kernel is not important anymore, as it is uniquely determined up to almost everywhere equality by the disintegration property.
As a corollary of the disintegration theorem for kernels, if $\mathcal{Z}$ is a standard Borel space then every finite measure on $\mathcal{Y}\times\mathcal{Z}$ admits a disintegration.
The proof of Theorem 3.2 goes through the construction of densities for finite kernels, and in particular it also gives a formalization of the Radon-Nikodym theorem for kernels. In order to state that theorem, we need to introduce the definition of a kernel with a density with respect to another kernel. First, for $\mu$ a measure on $\mathcal{Y}$ and $f:\mathcal{Y}\to\mathbb{R}_{+,\infty}$ a measurable function, we define a measure denoted by $f\cdot\mu$, the measure with density $f$ with respect to $\mu$, by $(f\cdot\mu)(B)=\int_{B}f(y)d\mu$ for all measurable sets $B$ of $\mathcal{Y}$ (Measure.withDensity). For $f:\mathcal{X}\to\mathcal{Y}\to\mathbb{R}_{+,\infty}$ a jointly measurable function (that is, a function such that the uncurried version is measurable on the product $\mathcal{X}\times\mathcal{Y}$) and $\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$ an s-finite kernel, we define the kernel $f\cdot\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$ (Kernel.withDensity) such that
$(f\cdot\eta)(x)(B)=\int_{y\in B}f(x)(y)d\eta(x)\>.$
That is, for all $x$, $(f\cdot\eta)(x)$ is defined as the measure $f(x)\cdot\eta(x)$ on $\mathcal{Y}$, and we can prove that thanks to the joint measurability of $f$ and the s-finiteness of $\eta$, $f\cdot\eta$ is indeed a kernel. In Mathlib we extend the definitions to functions that may not be measurable, by using the default value 0 is $f$ is not jointly measurable.
Recall that a measure $\mu$ is absolutely continuous with respect to a measure $\nu$ if for all measurable sets $B$, $\nu(B)=0$ implies $\mu(B)=0$, and is singular with respect to $\nu$ if there exists a measurable set $A$ such that $\mu(A)=0$ and $\nu(\mathcal{Y}\setminus A)=0$.
With those definitions, we can state the Radon-Nikodym theorem for kernels.
@@
Suppose that either $\mathcal{X}$ is countable or $\mathcal{Y}$ has a countably generated $\sigma$-algebra. Let $\kappa,\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$ be finite kernels. Then there exists a jointly measurable function $\frac{d\kappa}{d\eta}:\mathcal{X}\to\mathcal{Y}\to\mathbb{R}_{+,\infty}$ and a finite kernel $\kappa_{\perp\eta}:\mathcal{X}\rightsquigarrow\mathcal{Y}$ such that for every $x\in\mathcal{X}$, $\kappa_{\perp\eta}(x)$ is singular with respect to $\eta(x)$ and $\kappa=\frac{d\kappa}{d\eta}\cdot\eta+\kappa_{\perp\eta}$ .
@@ 3.1 Proof of the disintegration theorem
The main result from which both the disintegration theorem and the Radon-Nikodym theorem are derived is the existence of densities for finite kernels. Note that there is an unfortunate clash of terminology here: what we call density in this section is not a function such that $\kappa=d\cdot\eta$ for some kernel $\eta$ as in Kernel.withDensity. It has similar properties (integrating the density with respect to a kernel gives another kernel), but its type is different. For two kernels $\kappa,\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$, we write $\kappa\leq\eta$ if for all $x\in\mathcal{X}$ and measurable set $B$ of $\mathcal{Y}$, $\kappa(x)(B)\leq\eta(x)(B)$.
@@
Let $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}\times\mathcal{Z}$ and $\nu:\mathcal{X}\rightsquigarrow\mathcal{Y}$ be two finite kernels with $\kappa_{\mathrm{fst}}\leq\nu$, where $\mathcal{Y}$ has a countably generated $\sigma$-algebra. Then there exists a function $d_{\kappa,\nu}:\mathcal{X}\to\mathcal{Y}\to\mathcal{P}(\mathcal{Z})\to\mathbb{R}$ jointly measurable on $\mathcal{X}\times\mathcal{Y}$ such that for all $x\in\mathcal{X}$ and all measurable sets $B$ of $\mathcal{Y}$ and $C$ of $\mathcal{Z}$, $\int_{y\in B}d_{\kappa,\nu}(x)(y)(C)\,d\nu(x)=\kappa(x)(B\times C)$.
We added a definition for the density, Kernel.density, and its main property is stated in the following lemma.
[…]
∫⁻ x in B, ENNReal.ofReal (density κ ν a x C) ∂(ν a) = κ a (B ×ˢ C)
@@
Using the fact that a standard Borel space embeds measurably into $\mathbb{R}$ (embeddingReal), we can reduce the proof of Theorem 3.2 to the case where $\mathcal{Z}=\mathbb{R}$: for a general standard Borel space, we map the kernel to one with values in $\mathbb{R}$, disintegrate it, and then pull back the conditional kernel to the original space. Thus from a finite kernel $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}\times\mathbb{R}$, we want to build a kernel $\eta:\mathcal{X}\times\mathcal{Y}\rightsquigarrow\mathbb{R}$ such that $\kappa=\kappa_{\mathrm{fst}}\otimes_{k}\eta$. Since $\eta$ takes values in the measures on $\mathbb{R}$, we can build them from their cumulative distribution functions. A cumulative distribution function (CDF) is a nondecreasing right-continuous function $F:\mathbb{R}\to\mathbb{R}$ with limit 0 at -$\infty$ and limit 1 at $\infty$, and characterizes a probability measure on $\mathbb{R}$. Nondecreasing right-continuous functions $\mathbb{R}\to\mathbb{R}$ are bundled in Mathlib in a type StieltjesFunction and the corresponding measure can be obtained with StieltjesFunction.measure.
@@
If $\mathcal{X}$ is countable, we can provide for each $x\in\mathcal{X}$ a measurable function $f:\mathcal{Y}\to\mathbb{Q}\to\mathbb{R}$ and proceed as above to obtain a kernel $\mathcal{Y}\rightsquigarrow\mathbb{R}$. Since $\mathcal{X}$ is countable, measurability is not an issue and we can put those kernels together into a kernel $\mathcal{X}\times\mathcal{Y}\rightsquigarrow\mathbb{R}$. If on the other hand $\mathcal{X}$ is not countable, we cannot proceed separately for each $x\in\mathcal{X}$ and have to build a function which is measurable on the product. We are able to do so if $\mathcal{Y}$ has a countably generated $\sigma$-algebra thanks to the density provided by Theorem 3.4. The density $d_{\kappa,\nu}:\mathcal{X}\to\mathcal{Y}\to\mathcal{P}(\mathbb{R})\to\mathbb{R}$ is jointly measurable on $\mathcal{X}\times\mathcal{Y}$ and such that for all $x\in\mathcal{X}$ and all measurable sets $B$ of $\mathcal{Y}$ and all $q\in\mathbb{Q}$, $\int_{y\in B}d_{\kappa,\nu}(x)(y)((-\infty,q])\,d\nu(x)=\kappa(x)(B\times(-\infty,q])$. It thus satisfies the integral condition, and we are able to check that the other conditions are also fulfilled.
The countability of $\mathbb{Q}$ is important when proving the right-continuity. We want to prove that for almost all $p\in\mathcal{X}\times\mathcal{Y}$, $d_{\kappa,\nu}(p)$ is right-continuous on $\mathbb{Q}$ . Since $\mathbb{Q}$ is countable, it is enough to prove that for all $q\in\mathbb{Q}$, for almost all $p$, $d_{\kappa,\nu}(p)$ is right-continuous at $q$. To prove the almost everywhere equality of the limit on the right and the value of the function at a point $q\in\mathbb{Q}$, it suffices to prove that the integrals of those functions are equal on any measurable set $A$ of $\mathcal{X}\times\mathcal{Y}$. And since we know the integrals of $d_{\kappa,\nu}$, we can prove the equality of integrals.
@@
From $\kappa,\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$ two finite kernels we want to build a jointly measurable function $\frac{d\kappa}{d\eta}:\mathcal{X}\to\mathcal{Y}\to\mathbb{R}_{+,\infty}$ and a finite kernel $\kappa_{\perp\eta}:\mathcal{X}\rightsquigarrow\mathcal{Y}$ such that for every $x\in\mathcal{X}$, $\kappa_{\perp\eta}(x)$ is singular with respect to $\eta(x)$ and $\kappa=\frac{d\kappa}{d\eta}\cdot\eta+\kappa_{\perp\eta}$ .
@@
Let $\kappa,\eta:\mathcal{X}\rightsquigarrow\mathcal{Y}$ be two finite kernels with $\kappa\leq\eta$, in which $\mathcal{Y}$ has a countably generated $\sigma$-algebra. Then there exists a function $d_{\kappa,\eta}:\mathcal{X}\to\mathcal{Y}\to\mathbb{R}$ jointly measurable on $\mathcal{X}\times\mathcal{Y}$ such that for all $x\in\mathcal{X}$ and all measurable sets $B$ of $\mathcal{Y}$, $\int_{y\in B}d_{\kappa,\eta}(x)(y)\,d\eta(x)=\kappa(x)(B)$.
We use the corollary to define the Radon-Nikodym derivative of $\kappa$ with respect to $\eta$ in two steps. First, we define an auxiliary function rnDerivAux which is defined as the density $d_{\kappa,\eta}$ of the Corollary when $\mathcal{Y}$ has a countably generated $\sigma$-algebra and as the element-wise Radon-Nikodym derivative when $\mathcal{X}$ is countable.
[…]
else haveI := hCoC.countableOrCountablyGenerated.resolve_left hC
density (map κ (fun a (a, ()))) η x y univ
In this auxiliary definition, density is the function from Theorem 3.4 and density (map κ (fun a (a, ()))) η x y univ is the function from Corollary 3.5 (which in the code is defined even if we don’t have $\kappa\leq\eta$ but has meaningful properties only in that case).
In the Radon-Nikodym Theorem 3.3, we do not have the condition $\kappa\leq\eta$. We thus apply rnDerivAux to $\kappa$ and $\kappa+\eta$, which satisfy that condition. The Radon-Nikodym derivative $\frac{d\kappa}{d\eta}$ (Kernel.rnDeriv) is defined by the ratio $\frac{d_{\kappa,\kappa+\eta}}{1-d_{\kappa,\kappa+\eta}}$.
[…]
ENNReal.ofReal (rnDerivAux κ + η) x y) / ENNReal.ofReal (1 - rnDerivAux κ + η) x y)
We also use rnDerivAux to define the singular part (Kernel.rnSingularPart) $\kappa_{\perp\eta}$ of $\kappa$ with respect to $\eta$, as $\kappa_{\perp\eta}=(d_{\kappa,\kappa+\eta}-(1-d_{\kappa,\kappa+\eta})\frac{d\kappa}{d\eta})\cdot(\kappa+\eta)$. With the conventions of Lean regarding division by zero, the term $(1-d_{\kappa,\kappa+\eta}(x,y))\frac{d\kappa}{d\eta}(x,y)=(1-d_{\kappa,\kappa+\eta}(x,y))\frac{d_{\kappa,\kappa+\eta}(x,y)}{1-d_{\kappa,\kappa+\eta}(x,y)}$ is equal to 0 when $d_{\kappa,\kappa+\eta}(x,y)=1$ and to $d_{\kappa,\kappa+\eta}(x,y)$ otherwise.
[…]
withDensity + η) (fun x y Real.toNNReal (rnDerivAux κ + η) x y)
- Real.toNNReal (1 - rnDerivAux κ + η) x y) * rnDeriv κ η x y)
It is then not hard to check that $\frac{d\kappa}{d\eta}$ and $\kappa_{\perp\eta}$ satisfy the required properties when the kernels are finite. Thanks to the Radon-Nikodym Theorem 3.3 which is now proved, the condition $\kappa\leq\eta$ in Corollary 3.5 can be relaxed to absolute continuity of $\kappa$ with respect to $\eta$ (that is, absolute continuity of $\kappa(x)$ with respect to $\eta(x)$ for all $x$).
@@
We now give a sketch of the proof of Theorem 3.4. Our proof follows [kallenberg2021, Theorem 9.27]. Recall that $\kappa:\mathcal{X}\rightsquigarrow\mathcal{Y}\times\mathcal{Z}$ and $\nu:\mathcal{X}\rightsquigarrow\mathcal{Y}$ are two finite kernels with $\kappa_{\mathrm{fst}}\leq\nu$, where $\mathcal{Y}$ has a countably generated $\sigma$-algebra. If we were interested in a single $x\in\mathcal{X}$, we could use the Radon-Nikodym theorem for measures to obtain a density. For $x\in\mathcal{X}$, $y\in\mathcal{Y}$ and $C$ a measurable set of $\mathcal{Z}$, let $\kappa(x)_{|\mathcal{Y}\times C}$ be the restriction of the measure $\kappa(x)$ to the set $\mathcal{Y}\times C$. Let then $(\kappa(x)_{|\mathcal{Y}\times C})_{\mathrm{fst}}$ be the marginal of that measure on $\mathcal{Y}$. Then we could define the density as $d^{\prime}_{\kappa,\nu}(x,y)(C)=\frac{d(\kappa(x)_{|\mathcal{Y}\times C})_{\mathrm{fst}}}{d\nu(x)}(y)$ and it would satisfy the integral condition and be measurable in $y$. The problem is to obtain a density that is jointly measurable on $\mathcal{X}\times\mathcal{Y}$.